### v2

# On the Theoretical Limitations of Embedding-Based Retrieval

Abstract
Vector embeddings have been tasked with an ever-increasing set of retrieval tasks over the years, with a nascent rise in using them for reasoning, instruction-following, coding, and more. These new benchmarks push embeddings to work for any query and any notion of relevance that could be given. While prior works have pointed out theoretical limitations of vector embeddings, there is a common assumption that these difficulties are exclusively due to unrealistic queries, and those that are not can be overcome with better training data and larger models. In this work, we demonstrate that we may encounter these theoretical limitations in realistic settings with extremely simple queries. We connect known results in learning theory, showing that the number of top-$k$ subsets of documents capable of being returned as the result of some query is limited by the dimension of the embedding. We empirically show that this holds true even if we directly optimize on the test set with free parameterized embeddings. We then create a realistic dataset called LIMIT that stress tests embedding models based on these theoretical results, and observe that even state-of-the-art models fail on this dataset despite the simple nature of the task. Our work shows the limits of embedding models under the existing single vector paradigm and calls for future research to develop new techniques that can resolve this fundamental limitation.
1 Introduction
Over the last two decades, information retrieval (IR) has moved from models dominated by sparse techniques (such as BM25 Robertson et al. (1995)) to those that use neural language models (LM) as their backbones (Lee et al., 2019; Craswell et al., 2020; Izacard et al., 2021; Wang et al., 2022). These neural models are predominantly used in a single vector capacity, where they output a single embedding representing the entire input (also known as dense retrieval). These embedding models are capable of generalizing to new retrieval datasets and have been tasked with solving increasingly complicated retrieval problems (Thakur et al., 2021; Enevoldsen et al., 2025; Lee et al., 2025).
In recent years this has been pushed even further with the rise of instruction-following retrieval benchmarks, where models are asked to represent any relevance definition for any query (Weller et al., 2025a; c; Song et al., 2025; Xiao et al., 2024; Su et al., 2024). For example, the QUEST dataset (Malaviya et al., 2023) uses logical operators to combine different concepts, studying the difficulty of retrieval for complex queries (e.g., “Moths or Insects or Arthropods of Guadeloupe”). On the other hand, datasets such as BRIGHT (Su et al., 2024) explore the challenges arising from different definitions of relevance by defining relevance in ways that require reasoning. One subtask includes reasoning over a given Leetcode problem (the query) to find other Leetcode problems that share a subtask (e.g. others problems using dynamic programming). Although models cannot solve these benchmarks yet, the community has proposed these problems in order to push the boundaries of what dense retrievers are capable of—which is now implicitly every task that could be defined.
Rather than proposing empirical benchmarks to gauge what embedding models can achieve, we seek to understand at a more fundamental level what the limitations are. Since embedding models use vector representations in geometric space, there exist well-studied fields of mathematical research (Papadimitriou & Sipser, 1982) that could be used to analyze these representations.
Our work aims to bridge this gap, connecting known theoretical results in linear algebra with modern advancements in neural information retrieval. We draw upon research in high-dimensional geometry to provide a lower bound on the embedding dimension needed to represent a given combination of relevant documents and queries. Specifically, we show that for a given embedding dimension $d$ there exist top-$k$ combinations of documents that cannot be returned—no matter the query—highlighting a theoretical and fundamental limit to embedding models.
Figure 1: A depiction of the LIMIT dataset creation process, based on theoretical limitations. We test all combinations of relevance for $N$ documents (i.e. in the figure, all combinations of relevance for three documents with two relevant documents per query) and instantiate it using a simple mapping.
To show that this theoretical limit is true for any retrieval model or training dataset, we test a setting where the vectors themselves are directly optimized with the test data. This allows us to empirically show how the embedding dimension enables the solving of retrieval tasks. We find that there exists a crucial point for each embedding dimension ($d$) where the number of documents is too large for the embedding dimension to encode all combinations. We then gather these crucial points for a variety of $d$ and show that this relationship can be modeled empirically with a polynomial function.
We also go one step further and construct a realistic but simple dataset based on these theoretical limitations (called LIMIT). [^1] Data and code are available at https://github.com/google-deepmind/limit Despite the simplicity of the task (e.g., who likes Apples? and Jon likes Apples, ...), we find it is very difficult for even state-of-the-art embedding models (Lee et al., 2025; Zhang et al., 2025) on MTEB (Enevoldsen et al., 2025), and practically impossible for models with small embedding dimensions using standard optimization techniques.
Overall, our work contributes: (1) a theoretical basis for the fundamental limitations of embedding models, (2) a best-case empirical analysis showing that this proof holds for any dataset instantiation (by free embedding optimization), and (3) a simple real-world natural language instantiation called LIMIT that even state-of-the-art embedding models cannot solve.
These results imply interesting findings for the community: on one hand we see neural embedding models becoming immensely successful. However, academic benchmarks test only a small amount of the queries that could be issued (and these queries are often overfitted to), hiding these limitations. Our work shows that as the tasks given to embedding models require returning ever-increasing combinations of top-$k$ relevant documents (e.g., through instructions connecting previously unrelated documents with logical operators), we will reach a limit of combinations they can represent.
Thus, the community should be aware of these limitations, both when creating evals and also by using alternate architectures—such as cross-encoders / multi-vector / more expressive similarity functions —when trying to handle the full range of instruction queries, i.e. any query and relevance definition.
2 Related Work
2.1 Neural Embedding Models
There has been immense progress on embedding models in recent years (Lee et al., 2019; Craswell et al., 2020; BehnamGhader et al., 2024), moving from simple web search (text-only) to advanced instruction-following and multi-modal representations. These models generally followed advancements in language models, such as pre-trained LMs (Hoffmann et al., 2022), multi-modal LMs (Li et al., 2024; Team, 2024), and advancements in instruction-following (Zhou et al., 2023; Ouyang et al., 2022). Some of the prominent examples in retrieval include CoPali (Faysse et al., 2024) and DSE (Ma et al., 2024) which focus on multimodal embeddings, Instructor (Su et al., 2022) and FollowIR (Weller et al., 2024a) for instruction following, and GritLM (Muennighoff et al., 2024) and Gemini Embeddings (Lee et al., 2025) for pre-trained LMs turned embedders.
Our work, though focused solely on textual representations for simplicity, applies to all modalities of single vector embeddings for any domain of dataset. As the space of things to represent grows (through instructions or multi-modality) they will increasingly run into these theoretical limitations.
2.2 Empirical tasks pushing the limits of dense retrieval
Retrieval models have been pushed beyond their initial use cases to handle a broad variety of areas. Notable works include efforts to represent a wide group of domains (Thakur et al., 2021; Lee et al., 2024), a diverse set of instructions (Weller et al., 2024a; Zhou et al., 2024; Oh et al., 2024; Weller et al., 2025b), and to handle reasoning over the queries (Xiao et al., 2024; Su et al., 2024). This has pushed the focus of embedding models from basic keyword matching to embeddings that can represent the full semantic meaning of language. As such, it is more common than ever to connect what were previously unrelated documents into the top-$k$ relevant set, [^2] You can imagine an easy way to connect any two documents merely by using logical operators, i.e. X and Y. increasing the number of combinations that models must be able to represent. This has motivated our interest in understanding the limits of what embeddings can represent, as current work expects it to handle every task.
Previous work has explored empirically the limits of models: Reimers & Gurevych (2020) showed that smaller dimension embedding models have more false positives, especially with larger-scale corpora. Ormazabal et al. (2019) showed the empirical limitations of models in the cross-lingual setting and Yin & Shen (2018) showed how embedding dimensions relate to the bias-variance tradeoff. In contrast, our work provides a theoretical connection between the embedding dimension and the top-k sets it can retrieve, while also showing empirical limitations.
2.3 Theoretical Limits of Vectors in Geometric Space
Understanding and finding nearest neighbors in semantic space has a long history in mathematics research, with early work such as the Voronoi diagram being studied as far back as 1644 and formalized in 1908 (Voronoi, 1908). The order-$k$ version of the Voronoi diagram (i.e. the Voronoi diagram partitioning the space into regions based on their closest $k$ points) is obviously connected to information retrieval and has been studied for many years (Clarkson, 1988). The number of such regions is equal to the number of unique retrieval sets of size $k$, however this quantity is notoriously difficult to bound tightly (Bohler et al., 2015; Lee, 1982; Chen et al., 2023).
We approach this problem from a different angle, asking not how many $k$-subsets a given configuration realizes, but rather what embedding dimension is necessary to realize all $k$-subsets with a guaranteed score margin. By applying a classical sphere-packing volume argument (Vershynin, 2018; Conway et al., 1999), we obtain a lower bound on the embedding dimension in terms of $n$, $k$, and the margin $\gamma$. Our result is conceptually related to the Johnson–Lindenstrauss lemma (Johnson et al., 1984), which gives a sufficient dimension to preserve pairwise distances among $n$ points; in contrast, our bound gives a necessary dimension to realize all retrieval sets with a margin. The role of the margin in controlling the complexity of realizable configurations also parallels classical results in statistical learning theory, including the fat-shattering dimension (Kearns & Schapire, 1994) and margin-based generalization bounds for linear classifiers (Bartlett, 2002; Vapnik, 1998), where larger margins similarly constrain the capacity of the hypothesis class.
3 Representational Capacity of Vector Embeddings
In this section we formally define the minimum embedding dimension required to satisfy a given retrieval objective, and draw on classical sphere-packing results from high-dimensional geometry to establish a lower bound. We note that this will be an extreme lower bound, as practical models have to deal with other constraints such as learning through gradient descent and using LM tokenization.
Setup.
Let $v_{1},\dots,v_{n}\in\mathbb{R}^{d}$ be unit [^3] For simplicity, as nearly all SoTA retrieval models use unit vectors. document vectors, and let queries be unit vectors $u\in\mathbb{R}^{d}$. Fix $\gamma>0$. A $k$-subset $S\subseteq[n]$ is realized with margin $\gamma$ if there exists a unit query $u_{S}$ such that
$\min_{i\in S}\langle u_{S},v_{i}\rangle\ \geq\ \max_{j\notin S}\langle u_{S},v_{j}\rangle\ +\ 2\gamma.$ (1)
Since $\langle u,v_{i}\rangle\in[-1,1]$ for unit vectors, any score gap is at most $2$, hence equation 1 is feasible only for $0<\gamma\leq 1$. Throughout, $\log$ denotes the natural logarithm.
Theorem 1 (Dimension lower bound).
Assume $1\leq k<n$ and that every $k$-subset $S\subseteq[n]$ is realized with margin $\gamma$ as in equation 1. Then
$\binom{n}{k}\ \leq\ \Bigl(1+\frac{1}{\gamma}\Bigr)^{d},\qquad\text{hence}\qquad d\ \geq\ \frac{\log\binom{n}{k}}{\log\!\bigl(1+1/\gamma\bigr)}.$ (2)
Proof.
Fix two distinct $k$-subsets $S\neq T$ and choose $i\in S\setminus T$ and $j\in T\setminus S$. Applying equation 1 to $S$ and to $T$ gives
$\langle u_{S},v_{i}-v_{j}\rangle\geq 2\gamma,\qquad\langle u_{T},v_{j}-v_{i}\rangle\geq 2\gamma.$
Adding yields $\langle u_{S}-u_{T},v_{i}-v_{j}\rangle\geq 4\gamma$. By Cauchy–Schwarz and (universally, for any unit vectors) $\|v_{i}-v_{j}\|\leq\|v_{i}\|+\|v_{j}\|=2$, we obtain $\|u_{S}-u_{T}\|\geq 2\gamma$. Thus the $M=\binom{n}{k}$ unit queries $\{u_{S}\}$ are pairwise $2\gamma$-separated, so the open balls $B(u_{S},\gamma)$ are disjoint. Moreover, since $\|u_{S}\|=1$, each $B(u_{S},\gamma)\subseteq B(0,1+\gamma)$, and therefore
$M\cdot\mathrm{vol}\!\bigl(B_{d}(\gamma)\bigr)\ \leq\ \mathrm{vol}\!\bigl(B_{d}(1+\gamma)\bigr).$
Using $\mathrm{vol}(B_{d}(r))=C_{d}\,r^{d}$ for a constant $C_{d}$ depending only on $d$ (which cancels), we get $M\gamma^{d}\leq(1+\gamma)^{d}$, i.e. $\binom{n}{k}\leq\bigl((1+\gamma)/\gamma\bigr)^{d}=\bigl(1+1/\gamma\bigr)^{d}$. Rearranging yields equation 2. ∎
Corpus size $n$ $k=2$ $k=10$ $k=100$ $k=1000$
$10^{2}$ 4 13 trivial
$10^{3}$ 6 23 135 trivial
$10^{4}$ 8 33 233 1354
$10^{5}$ 10 42 329 2334
$10^{6}$ 12 52 425 3296
$10^{7}$ 14 61 521 4257
$10^{8}$ 16 71 617 5217
$10^{9}$ 17 81 713 6177
$10^{10}$ 19 90 809 7137
$10^{11}$ 21 100 905 8098
Table 1: Lower bounds for embedding dimension from Theorem 1 for $\gamma=0.1$. When $n$ and $k$ are both 1000, the result is trivial because $\binom{1000}{1000}=1$ (there is only one $k$-subset, hence no “irrelevant” items to separate). We see for large $k$ and $n$ values these dimension requirements are already greater than those currently used for web-scale search. If these numbers are inflated by a small multiple due to constraints on gradient learning or other LM-based constraints (e.g. tokenization, generalization) these bounds are outside of any reasonable embedding dimension.
3.0.1 Implications
Numerical instantiation
We can illustrate the effects of this lower bound using $\gamma=0.1$ (score gap $2\gamma=0.2$), which is approximately standard for models based on empirical usage. Thus, equation 2 becomes $d\geq\left\lceil\log\binom{n}{k}/\log 11\right\rceil$, with a table for various $k$ and $n$ values in Table 1.
For $n\gg k$, $\log\binom{n}{k}\approx k\log(en/k)$, so equation 2 forces
$d\ =\ \Omega\!\left(\frac{k\log(en/k)}{\log(1+1/\gamma)}\right).$
A stricter margin requirement (larger $\gamma$) demands higher dimension, since $\log(1+1/\gamma)$ decreases with $\gamma$ (feasibility requires $\gamma\leq 1$, so the denominator is at least $\log 2$).
Consequences
Due to space and speed requirements, most embeddings used for web-scale search are quantized or truncated (e.g. through Matryoshka embeddings (Kusupati et al., 2022)) to less than 1k dimensions, while the largest embeddings used in research are around 4096 (Zhang et al., 2025). We see that even with a moderate margin, which is needed to handle noise from messy data or quantization, the lower bounds in Table 1 can already be larger than what is used in practice.
Additional constraints on real-world models (such as needing to generalize, learn from gradient descent, and use natural language and tokenization) will make the dimension required in practice much higher. As Table 1 shows, even a small multiple of this lower bound would make the embedding dimension requirement infeasible. This multiple seems well-founded, as we will show in the next section from the best-case optimization setting (e.g. free embeddings).
4 Empirical Connection: Best Case Optimization
Having established a theoretical limitation of embedding models based on their embedding dimension $d$, we seek to show that this holds empirically also.
To show the strongest optimization case possible, we design experiments where the vectors themselves are directly optimizable with gradient descent. [^4] This could also be viewed as an embedding model where each query/doc are a separate vector via lookup. We call this “free embedding” optimization, as the embeddings are free to be optimized and not constrained by natural language, which imposes constraints on any realistic embedding model. Thus, this shows whether it is feasible for any embedding model to solve this problem: if the free embedding optimization cannot solve the problem, real retrieval models will not be able to either. It is also worth noting that we do this by directly optimizing the embeddings over the target qrel matrix (test set). This will not generalize to a new dataset, but is done to show the highest performance that could possibly occur.
Experimental Settings
[…]
We focus on relatively small sizes for $n$, $k$, and $d$ due to the combinatorial explosion of combinations with larger document values (i.e. 50k docs with top-$k$ of 100 gives 7.7e+311 combinations, which would be equivalent to the number of query vectors of dimension $d$ in that free embedding experiment). We use $k=2$ and increase $n$ by one for each $d$ value until it breaks. We fit a polynomial regression line to the data so we can model and extrapolate results outwards.
Figure 2: The critical-n value where the dimensionality is too small to successfully represent all the top-2 combinations. We plot the trend line as a polynomial function.
Results
Figure 2 shows that the curve fits a 3rd degree polynomial curve, with formula $y=-10.5322+4.0309d+0.0520d^{2}+0.0037d^{3}$ ($r^{2}$=0.999). Extrapolating this curve outward gives the critical-n values (for embedding size): 500k (512), 1.7m (768), 4m (1024), 107m (3072), 250m (4096). We note that this is the best case: a real embedding model cannot directly optimize the query and document vectors to match the test qrel matrix (and is constrained by factors such as “modeling natural language”). The results also show that the lower bounds in the previous section are a gross underestimate of real-world performance, as Table 1 shows a lower bound of 4 for $n=100$ whereas we see the free embeddings needing $d>18$ (e.g. a 4.5 multiplier even in the no-generalization or natural language case). Overall, these numbers already show that for web-scale search, even the largest embedding dimensions with ideal test-set optimization are not enough to model all combinations.
5 Empirical Connection: Real-World Datasets
The free embedding experiments provide empirical evidence that our theoretical results hold true. However, they still are abstract - what does this mean for real embedding models? In this section we (1) draw connections from this theory to existing datasets and (2) create a trivially simple yet extremely difficult retrieval task for existing SOTA models.
5.1 Connection to Existing Datasets
Existing retrieval datasets typically use a static evaluation set with limited numbers of queries, as relevance annotation is expensive to do for each query. This means practically that the space of queries used for evaluation is a very small sample of the number of potential queries. For example, the QUEST dataset (Malaviya et al., 2023) has 325k documents and queries with 20 relevant documents per query, with a total of 3357 queries. The number of unique top-20 document sets that could be returned with the QUEST corpus would be $\binom{325k}{20}$ which is equal to 7.1e+91 (larger than the estimate of atoms in the observable universe, $10^{82}$). Thus, the 3k queries in QUEST can only cover an infinitesimally small part of the qrel combination space.
Although it is not possible to instantiate all combinations when using large-scale corpora, search evaluation datasets are a proxy for what any user would ask for and ideally would be designed to test many combinations, as users will do. In many cases, developers of new evaluations simply choose to use fewer queries due to cost or computational expense of evaluation. For example, QUEST’s query “Novels from 1849 or George Sand novels” combines two categories of novels with the “OR” operator – one could instantiate new queries to relate concepts through OR’ing other categories together. Similarly, with the rise of search agents, we see greater usage of hyper-specific queries: BrowseComp (Wei et al., 2025) has 5+ conditions per query, including range operators. With these tools, it is possible to sub-select any top-$k$ relevant set with the right operators if the documents are sufficiently expressive (i.e. non-trivial). Thus, that existing datasets choose to only instantiate some of these combinations is mainly for practical reasons and not because of a lack of existence.
In contrast to these previous works, we seek to build a dataset that evaluates all combinations of top-$k$ sets for a small number of documents. Rather than using difficult query operators like QUEST, BrowseComp, etc. (which are already difficult for reasons outside of the qrel matrix) we choose very simple queries and documents to highlight the difficulty of representing all top-$k$ sets themselves.
5.2 The LIMIT Dataset
Dataset Construction
In order to have a natural language version of this dataset, we need some way to map combinations of documents into something that could be retrieved with a query. One simple [^7] This is just one way, designed to be realistic and simple. However, our framework allows for any way of instantiation – not stuck to this arbitrary natural language design. way to do this is to create a synthetic version with latent variables for queries and documents and then instantiate it with natural language. For this mapping, we choose to use attributes that someone could like (i.e. Jon likes Hawaiian pizza, sports cars, etc. ) as they are plentiful and don’t present issues w.r.t. other items: one can like Hawaiian pizza but dislike pepperoni, all preferences are valid. We then enforce two constraints for realism: (1) users shouldn’t have too many attributes, thus keeping the documents short (less than 50 per user) and (2) each query should only ask for one item to keep the task simple (i.e. “who likes X”). We gather a list of attributes a person could like through prompting Gemini 2.5 Pro. We then clean it to a final 1850 items by iteratively asking it to remove duplicates/hypernyms, while also checking the top failures with BM25 to ensure no overlap.
We choose to use 50k documents in order to have a hard but relatively small corpus and 1000 queries to maintain statistical significance while still being fast to evaluate. For each query, we choose to use two relevant documents (i.e. kk=2), both for simplicity in instantiating and to mirror previous work (i.e. NQ, HotpotQA, etc. (Kwiatkowski et al., 2019; Yang et al., 2018)).
Figure 3: Scores on the LIMIT task. Despite the simplicity of the task we see that SOTA models struggle. We also see that the dimensionality of the model is a limiting factor and that as the dimension increases, so does performance. Even multi-vector models struggle. Lexical models like BM25 do very well due to their higher dimensionality. Stars indicate models trained with MRL.
Our last step is to choose a qrel matrix to instantiate these attributes. Although we could not prove the hardest qrel matrix definitively with theory, we intuit that our theoretical results imply that the more interconnected the qrel matrix is (e.g. dense with all combinations) the harder it would be for models to represent. Following this, we use the qrel matrix with the highest number of documents for which all combinations would be just above 1000 queries for a top-$k$ of 2 (46 docs, since $\binom{46}{2}$ is 1035, the smallest above 1k).
We then assign random natural language attributes to the queries, adding these attributes to their respective relevant documents (c.f. Figure 1). We give each document a random first and last name from open-source lists of names. Finally, we randomly sample new attributes for each document until all documents have the same number of attributes. As this setup has many more documents than those that are relevant to any query (46 relevant documents, 49.95k non-relevant to any query) we also create a “small” version with only the 46 documents that are relevant to one of the 1000 queries.
Models   We evaluate the state-of-the-art embedding models including GritLM (Muennighoff et al., 2024), Qwen 3 Embeddings (Zhang et al., 2025), Promptriever (Weller et al., 2024b), Gemini Embeddings (Lee et al., 2025), Snowflake’s Arctic Embed Large v2.0 (Yu et al., 2024), and E5-Mistral Instruct (Wang et al., 2022; 2023). These models range in embedding dimension (1024 to 4096) as well as in training style (instruction-based, hard negative optimized, etc.). We also evaluate three non-single vector models to show the distinction: BM25 (Robertson et al., 1995; Lù, 2024), gte-ModernColBERT (Chaffin, 2025a; Chaffin & Sourty, 2024), and a token-wise TF-IDF. [^8] This model turns each unique item into a token and then does TF-IDF. We build it to show that it gets 100% on all tasks (as it reverse engineers our dataset construction) and thus we do not include it in future charts.
We show results at the full embedding dimension and also with truncated embedding dimension (typically used with matryoshka learning, aka MRL (Kusupati et al., 2022)). For models not trained with MRL this will result in sub-par scores, thus, models trained with MRL are indicated with stars in the plots. However, as there are no LLMs with an embedding dimension smaller than 384, we include MRL for all models to small dimensions (32) to show the impact of embedding dimensionality.
Figure 4: Scores on the LIMIT small task (N=46) over embedding dimensions. Despite having just 46 documents, models struggle even with recall@10 and cannot solve the task even with recall@20.
Results   Figure 3 shows the results on the full LIMIT while Figure 4 shows the results on the small (46 document) version. The results are surprising - models severely struggle even though the task is trivially simple. For example, in the full setting models struggle to reach even 20% recall@100 and in the 46 document version models cannot solve the task even with recall@20.
Figure 5: Training on LIMIT train does not significantly help, indicating the issue is not domain shift. But models can solve it if they overfit to the test set.
We see that model performance depends crucially on the embedding dimensionality (better performance with bigger dimensions). Interestingly, models trained with more diverse instruction, such as Promptriever, perform better, perhaps because their training allows them to use more of their embedding space (compared to models which are trained with MRL and on a smaller range of tasks that can perhaps be consolidated into a smaller embedding manifold).
For alternative architectures, GTE-ModernColBERT does significantly better than single-vector models (although far from solving the task) while BM25 comes close to perfect scores. Both of these alterative architectures (sparse and multi-vector) offer various trade-offs, see section˜5.3 for analysis.
Is this Domain Shift?   Although our queries look similar to standard web search queries, we wondered whether there could be some domain shift causing the low performance. If so, we would expect that training on a training set of similar examples would significantly improve performance. On the other hand, if the task was intrinsically hard, training on the training set would provide little help whereas training on the test set would allow the model to overfit to those tokens (similar to the free embeddings).
To test this we take an off-the-shelf embedding model and train it on either the training set (created synthetically using non-test set attributes) or the official test set of LIMIT. We use lightonai/modernbert-embed-large (Chaffin, 2025c) and fine-tune it on these splits, using the full dataset for in batch negatives (excluding positives) using SentenceTransformers (Reimers & Gurevych, 2019). We show a range of dimensions by projecting the hidden layer down to the specified size during training (rather than using MRL).
Figure˜5 shows the model trained on the training set cannot solve the problem, although it does see very minor improvement from near zero recall@10 to up to 2.8 recall@10. The lack of performance gains when training in-domain indicate that poor performance is not due to domain shift. By training the model on the test set we see it can learn the task, overfitting on the tokens in the test queries. This aligns with our free embedding results, that it is possible to overfit to the $N=46$ version with only 12 dimensions. However, it is notable that the real models with 64 dimensions still cannot completely solve the task, implying real models perform significantly worse than the bounds shown in §4.
What about Non-Lexical Matches?   Our previous results show that lexical models greatly outperform their neural counterparts. However, this is not to imply that lexical models are a panacea - although they have higher dimensionality than single-vector models, they have other shortcomings.
We illustrate this by creating a version of LIMIT-small that replaces all items in the corpus with their synonyms, reducing the amount of lexical overlap. We ask Gemini 2.5 Pro to come up with synonyms that don’t match any other existing synonyms or original items, by using either scientific names, similar meanings, or if necessary hypernyms. This creates a mapping like “glasses” to “spectacles”, etc. [^9] We note that it doesn’t remove all lexical overlap due to items like “Scuba Diving” -¿ “Underwater Diving” We repeat the previous experiment on LIMIT-small (synonyms) and compare to LIMIT-small in Figure 6. We find that all models drop in performance as the task is now more difficult, but BM25 drops the most and now underperforms the neural models (e.g. BM25 drops more than 89% whereas Qwen3 embedding drops 38.9%). Thus, we can see that although lexical models have strengths like higher dimensionality, they are limited by their keyword-only matching ability. We expand more upon their strengths and weaknesses for instruction-following in Section 5.3.
Figure 6: Comparing scores on LIMIT small vs LIMIT small (synonym). Using synonyms makes the task harder so all models perform worse. However, the lexical model (BM25) drops nearly 90%, performing worse than most single-vector models on the synonym version of LIMIT. Thus, lexical models have weaknesses of their own (see Section 5.3 for more discussion) and are not a panacea.
Implications   Single-vector models are fundamentally limited by their embedding dimension. The LIMIT dataset is a particular instantiation, with very simple queries and documents, designed to highlight this property. This small version of LIMIT can be embedded in just 12 dimensions (as seen in the free embeddings experiments), yet all models fail to perform well, suggesting other architectural weaknesses. Irrespective of the architecture involved, however, our framework can scale the dataset’s difficulty to consistently demonstrate this fundamental limitation.
5.3 Alternatives to Embedding Models
Our previous results show both theoretically and empirically that embedding models cannot represent all combinations of documents in their top-$k$ sets, making them unable to represent and solve some retrieval tasks. As current embedding models have grown larger (e.g. up to 4096), this has helped reduce negative effects for smaller dataset sizes. However, with enough combinations of top-$k$ sets the dimensionality would have to increase to an infeasible size for non-toy datasets. Thus, although they are useful for first stage results, more expressive retriever architectures will be needed.
Cross-Encoders   Although not suitable for first stage retrieval at scale, they are already typically used to improve first stage results. Is LIMIT challenging for rerankers also? We evaluate a long context reranker, Gemini-2.5-Pro (Comanici et al., 2025) on the small setting as a comparison. We give Gemini all 46 documents and all 1000 queries at once, asking it to output the relevant documents for each query with one generation. We find that it can successfully solve (100%) all 1000 queries in one forward pass. This is in contrast to even the best embedding models with a recall@2 of less than 60% (Figure 4). Thus we can see that LIMIT is easy for state-of-the-art reranker models, which do not have the same limitations based on embedding dimension.
Multi-vector models   Multi-vector models are more expressive through the use of multiple vectors per sequence combined with the MaxSim operator (Khattab & Zaharia, 2020). These models show promise on the LIMIT dataset, with scores greatly above the single-vector models despite using a smaller backbone (ModernBERT, Warner et al. (2024)). However, these models are not generally used for instruction-following or reasoning-based tasks (see Chaffin (2025b) as one of the few that exist), leaving it an open question to how well multi-vector techniques will transfer to these tasks.
Sparse models   Sparse models (both lexical and neural) can be thought of as single vectors but with very high dimensionality. This dimensionality helps BM25 avoid the problems of the neural embedding models as seen in Figure 3. Since the $d$ of their vectors is high, they can scale to many more combinations than their dense vector counterparts. However, it is less clear how to apply sparse models to instruction-following and reasoning-based tasks where there is no lexical or even paraphrase-like overlap. We leave this direction (and hybrid sparse/dense solutions) to future work.
We note that all of these options have various trade-offs and none provide a clear path to solving this problem as-is. We leave it to future work to develop new techniques to mitigate these issues: perhaps through one of these alterative categories or through new ideas around single-vector models that can resolve the underlying issue (potentially through techniques such as hyperencoders (Killingback et al., 2025) or other future work on single vector architectures yet to be developed).
6 Conclusion
We introduce the LIMIT dataset, which highlights a fundamental limitation of embedding models. We provide a theoretical connection which shows that, for a fixed embedding dimension there will be some set of documents such that certain sets are unattainable as top-$k$ sets. We show these theoretical results hold empirically, through best case optimization of the vectors themselves, and make a practical connection to existing state-of-the-art models by creating a realistic and simple instantiation of the theory, called LIMIT, that these models cannot solve. Our results imply that the community should reconsider how instruction-based retrieval will impact future retrievers.
Limitations
Although our experiments provide theoretical insight for the most common type of embedding model (single vector) they do not hold necessarily for other architectures, such as multi-vector models. Although we showed initial empirical results with non-single vector models, we leave it to future work to extend our theoretical connections to these settings. We also did not show theoretical results for the setting where the user allows some mistakes, e.g. capturing only the majority of the combinations. We leave putting a bound on this scenario to future work and would invite the reader to examine works like Ben-David et al. (2002).
We have shown the theoretical connection that proves that some combinations cannot be represented by embedding models, however, we cannot prove apriori which types of combinations they will fail on. Thus, it is possible that there are some instruction-following or reasoning tasks they can solve perfectly, however, we do know that there exists some tasks that they will never be able to solve.
Acknowledgments
We thank Tanmaya Dabral, Zhongli Ding, Anthony Chen, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova for their helpful feedback. We thank Kiril Bangachev, Guy Bresler, Iliyas Noman, Yury Polyanskiy, Antonio Vergari, Adam Lopez, and Andreas Grivas for pointers to work on sign-rank.
Appendix A Relationship to Order-K Voronoi Regions
We also provide an explanation for how our results compare to Clarkson (1988) which put bounds on the number of regions in the order-$k$ Voronoi graph. The order-$k$ Voronoi graph is defined as the set of points having a particular set of $n$ points in $S$ as its $n$ nearest neighbors. This maps nicely to retrieval, as each order-$k$ region is equivalent to one retrieved set of top-$k$ results. Then the count of unique regions in the Voronoi graph is the total number of combinations that could be returned for those points. However, creating an empirical order-k Voronoi graph is computationally infeasible for $d$ > 3, and theoretically it is hard to bound tightly. Thus we use a different approach for showing the limitations of embedding models.
Appendix B Hyperparameter and Compute Details
Inference
We use the default length settings for evaluating models using the MTEB framework (Enevoldsen et al., 2025). As our dataset has relatively short documents (around 100 tokens), this does not cause an issue.
Training
For training on the LIMIT training and test set we use the SentenceTransformers library (Reimers & Gurevych, 2019) using the MultipleNegativesRankingLoss. We use a full dataset batch size and employ the no duplicates sampler to ensure that no in-batch negatives are duplicates of the positive docs. We use a learning rate of 5e-5. We train for 5 epochs and limit the training set slightly to the size of the test set (from 2.5k to 2k examples, matching test).
Compute
Inference and training for LIMIT is done with A100 GPUs on Google Colab Pro. The free embedding experiments are done mainly on H100 GPUs and TPU v5’s for larger size $N$ to accommodate higher VRAM for full-dataset batch vector optimization.
Appendix C Sigmoid Loss Function for Free Embeddings
In concurrent work by Bangachev et al. (2025), they show that for vision-language embedding models like CLIP (that more commonly use sigmoid loss functions) that the free-embedding experiments can be solved in fewer dimensions than in our setting (assuming no margin). As our results found the best performance with InfoNCE, which attempts to create the widest possible margin, this indicates that there are additional questions to resolve around learnability. We welcome further insight into this question both theoretically and empirically, as there exists widely disparate practices between the vision-language community (where sigmoid loss functions are often SOTA) and the text-only community (where sigmoid loss functions are almost never used due to worse performance).
This sigmoid learning is also closely related to other work, such as Grivas et al. (2024) and generally on the topic of work such as (Badreddine et al., 2025; Paul et al., 2021)
Appendix D Proof using Sign-Rank
In the initial version of this paper, we provided a theoretical bound without any margin requirement based on the qrel matrices sign rank. Although the proof is correct, the sign rank of the $\binom{n}{k}$ matrix has been established in previous work (Alon et al., 1985) and only depends on k. We include the proof connecting notions relevant for retrieval with the classic notion of sign-rank, however we emphasize that this will provide a weaker requirement on dimension as it assumes no margin.
D.1 Formalization
We consider a set of $m$ queries and $n$ documents with a ground-truth relevance matrix $A\in\{0,1\}^{m\times n}$, where $A_{ij}=1$ if and only if document $j$ is relevant to query ii. [^10] The matrix $A$ is often called the “qrels” (query relevance judgments) matrix in information retrieval. Vector embedding models map each query to a vector $u_{i}\in\mathbb{R}^{d}$ and each document to a vector $v_{j}\in\mathbb{R}^{d}$. Relevance is modeled by the dot product $u_{i}^{T}v_{j}$, with the goal that relevant documents should score higher than irrelevant ones.
Concatenating the vectors for queries in a matrix $U\in\mathbb{R}^{d\times m}$ and those for documents in a matrix $V\in\mathbb{R}^{d\times n}$, these dot products are the entries of the score matrix $B=U^{T}V$. The smallest embedding dimension $d$ that can realize a given score matrix is, by definition, the rank of $B$. Therefore, our goal is equivalent to finding the minimum rank of a score matrix $B$ that correctly orders documents according to the relevance specified in $A$, which we formalize in the following definition.
Definition 1.
Given a matrix $A\in\mathbb{R}^{m\times n}$, the row-wise order-preserving rank of $A$ is the smallest integer $d$ such that there exists a rank-$d$ matrix $B$ that preserves the relative order of entries in each row of $A$. We denote this as
[…]
In other words, if $A$ is a binary ground-truth relevance matrix, $\operatorname{rank}_{\text{rop}}A$ is the minimum dimension necessary for any vector embedding model to return relevant documents before irrelevant ones for all queries. Alternatively, we might require that the scores of relevant documents can be cleanly separated from those of irrelevant ones by a threshold.
Definition 2.
Given a binary matrix $A\in\{0,1\}^{m\times n}$:
  • The row-wise thresholdable rank of AA ($\operatorname{rank}_{\text{rt}}A$) is the minimum rank of a matrix $B$ for which there exist row-specific thresholds $\{\tau_{i}\}_{i=1}^{m}$ such that for all $i,j$, $B_{ij}>\tau_{i}$ if $A_{ij}=1$ and $B_{ij}<\tau_{i}$ if $A_{ij}=0$.
  • The globally thresholdable rank of AA ($\operatorname{rank}_{\text{gt}}A$) is the minimum rank of a matrix $B$ for which there exists a single threshold $\tau$ such that for all $i,j$, $B_{ij}>\tau$ if $A_{ij}=1$ and $B_{ij}<\tau$ if $A_{ij}=0$.
Remark 1.
This two-sided separation condition may be seen as slightly stronger than requiring $B_{ij}>\tau_{i}$ if and only if $A_{ij}=1$, however since there are only finitely many elements of $B_{ij}$ we could always perturb the latter threshold by a sufficient number such that the two-sided condition holds. [^11] Without loss of generality, we may assume the thresholds in the above definitions are not equal to any elements of $B$ since we could increase the threshold of $\tau$ by a sufficiently small $\epsilon$ to preserve the inequality.
D.2 Theoretical Bounds
For binary matrices, row-wise ordering/thresholding are equivalent notions of representation capacity.
Proposition 1.
For a binary matrix $A\in\{0,1\}^{m\times n}$, we have that $\operatorname{rank}_{\text{rop}}A=\operatorname{rank}_{\text{rt}}A$.
Proof.
($\leq$) Suppose $B$ and $\tau$ satisfy the row-wise thresholdable rank condition. Since $A$ is a binary matrix $A_{ij}>A_{ik}$ implies $A_{ij}=1$ and $A_{ik}=0$, thus $B_{ij}>\tau_{i}>B_{ik}$, and hence $B$ also satisfies the row-wise order-preserving condition.
($\geq$) Let $B$ satisfy the row-wise order-preserving condition, so $A_{ij}>A_{ik}$ implies $B_{ij}>B_{ik}$. For each row $i$, let $U_{i}=\{B_{ij}\mid A_{ij}=1\}$ and $L_{i}=\{B_{ij}\mid A_{ij}=0\}$. The row-wise order-preserving condition implies that every element of $U_{i}$ is greater than every element of $L_{i}$. We can therefore always find a threshold $\tau_{i}$ separating them (e.g. $\tau_{i}=(\max L_{i}+\min U_{i})/2$ if both are non-empty, trivial otherwise). Thus $B$ is also row-wise thresholdable to $A$. ∎
The notions we have described so far are closely related to the sign rank of a matrix, which we use in the rest of the paper to establish our main bounds.
Definition 3 (Sign Rank).
The sign rank of a matrix $M\in\{-1,1\}^{m\times n}$ is the smallest integer $d$ such that there exists a rank $d$ matrix $B\in\mathbb{R}^{m\times n}$ whose entries have the same sign as those of $M$, i.e.
[…]
In what follows, we use $\mathbf{1}_{n}$ to denote the $n$-dimensional vector of ones, and $\mathbf{1}_{m\times n}$ to denote an $m\times n$ matrix of ones.
Proposition 2.
Let $A\in\{0,1\}^{m\times n}$ be a binary matrix. Then $2A-\mathbf{1}_{m\times n}\in\{-1,1\}^{m\times n}$ and
$\operatorname{rank}_{\pm}(2A-\mathbf{1}_{m\times n})-1\leq\operatorname{rank}_{\text{rop}}A=\operatorname{rank}_{\text{rt}}A\leq\operatorname{rank}_{\text{gt}}A\leq\operatorname{rank}_{\pm}(2A-\mathbf{1}_{m\times n})$
Proof.
N.b. the equality was already shown in Proposition˜1. We prove each inequality separately.
1. $\operatorname{rank}_{\text{rt}}A\leq\operatorname{rank}_{\text{gt}}A$: True by definition, since any matrix satisfying the globally thresholdable condition trivially satisfies a row-wise thresholdable condition with the same threshold for each row.
2. $\operatorname{rank}_{\text{gt}}A\leq\operatorname{rank}_{\pm}(2A-\mathbf{1}_{m\times n})$: Let $B$ be any matrix whose entries have the same sign as $2A-\mathbf{1}_{m\times n}$, then
$B_{ij}>0\iff 2A_{ij}-1>0\iff A_{ij}=1.$
Thus $B$ satisfies the globally thresholdable condition with a threshold of 0.
3. $\operatorname{rank}_{\pm}(2A-\mathbf{1}_{m\times n})-1\leq\operatorname{rank}_{\text{rt}}A$: Suppose $B$ satisfies the row-wise thresholdable condition with minimal rank, so $\operatorname{rank}_{\text{rt}}A=\operatorname{rank}B$ and there exists $\tau\in\mathbb{R}^{m}$ such that $B_{ij}>\tau_{i}$ if $A_{ij}=1$ and $B_{ij}<\tau_{i}$ if $A_{ij}=0$. Then the entries of $B-\tau\mathbf{1}_{n}^{T}$ have the same sign as $2A-\mathbf{1}_{m\times n}$, since $(B-\tau\mathbf{1}_{n}^{T})_{ij}=B_{ij}-\tau_{i}$ and
$B_{ij}-\tau_{i}>0$ $\iff A_{ij}=1\iff 2A_{ij}-1>0,\text{ and}$ (3)
$B_{ij}-\tau_{i}<0$ $\iff A_{ij}=0\iff 2A_{ij}-1<0.$ (4)
Thus $\operatorname{rank}_{\pm}(2A-\mathbf{1}_{m\times n})\leq\operatorname{rank}(B-\tau\mathbf{1}_{n}^{T})\leq\operatorname{rank}(B)+\operatorname{rank}(\tau\mathbf{1}_{n}^{T})=\operatorname{rank}_{\text{rt}}A+1$.
Combining these gives the desired chain of inequalities. ∎
D.3 Consequences
In the context of a vector embedding model, this provides a lower and upper bound on the dimension of vectors required to exactly capture a given set of retrieval objectives, in the sense of row-wise ordering, row-wise thresholding, or global thresholding. In particular, given some binary relevance matrix $A\in\{0,1\}^{m\times n}$, we need at least $\operatorname{rank}_{\pm}(2A-\mathbf{1}_{m\times n})-1$ dimensions to capture the relationships in $A$ exactly, and can always accomplish this in at most $\operatorname{rank}_{\pm}(2A-\mathbf{1}_{m\times n})$ dimensions.
The cyclotomic polynomial construction presented in Alon et al. (1985) implies that any qrel matrix has sign-rank at most $2k$, where $k$ is the largest number of documents for a particular query. The construction results in unnormalized vectors, however this can be easily adapted to normalized vectors by using one additional dimension. In agreement with Theorem˜1, this construction requires infinite precision in general, and is thus not feasible in practice.
D.4 Correlation with MTEB
Figure 7: No obvious correlation between BEIR vs LIMIT.
BEIR (used in MTEB v1) (Thakur et al., 2021; Muennighoff et al., 2022) has frequently been cited as something that embedding models have overfit to (Weller et al., 2025c; Thakur et al., 2025). We compare performance on LIMIT to BEIR in Figure 7. We see that performance is generally not correlated and that smaller models (like Arctic Embed) do worse on both, likely due to embedding dimension and pre-trained model knowledge.
Appendix E LLM Usage
LLMs were not used for any paper writing, only for coding help and title brainstorming.
Appendix F Metrics Measuring Qrel Graph Density
We show two metrics that treat the qrel matrix as a graph and show that LIMIT has unique properties compared to standard IR datasets (Table 2). We call these metrics Graph Density and Average Query Strength and describe them below.
Graph Density
We use the qrel matrix to construct the graph, where nodes are documents and an edge exists between two documents if they are both relevant to at least one common query.
For a given graph $G=(V,E)$ with $V$ being the set of nodes and $E$ being the set of edges, the graph density is defined as the ratio of the number of edges in the graph to the maximum possible number of edges. For an undirected graph, the maximum possible number of edges is $\frac{|V|(|V|-1)}{2}$. Thus, the density $\rho$ is calculated as:
$\rho=\frac{|E|}{\frac{|V|(|V|-1)}{2}}=\frac{2|E|}{|V|(|V|-1)}$
This metric indicates how connected the graph is; a density of 1 signifies a complete graph (all possible edges exist), while a density close to 0 indicates a sparse graph. For a qrel dataset, the
Average Query Strength
In a query-query graph where nodes are queries and edges represent similarity between queries (e.g., Jaccard similarity of their relevant documents), the strength of a query node $i$, denoted $s_{i}$, is defined as the sum of the weights of all edges incident to it. If $w_{ij}$ is the weight of the edge between query $i$ and query $j$, and $N(i)$ is the set of neighbors of query $i$, then the strength is:
$s_{i}=\sum_{j\in N(i)}w_{ij}$
The Average Query Strength $\bar{s}$ is the mean of these strengths across all query nodes in the graph:
$\bar{s}=\frac{1}{|V_{Q}|}\sum_{i\in V_{Q}}s_{i}$
where $V_{Q}$ is the set of all query nodes in the graph. This metric provides an overall measure of how strongly connected queries are to each other on average within the dataset, based on their shared relevant documents.
Comparisons to other datasets
We compare with standard IR Datasets such as NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and SciFact (Wadden et al., 2020). We also show an instruction-following dataset, FollowIR Core17 (Weller et al., 2024a). For all datasets, we use the test set only. The results in Table 2 show that LIMIT has significantly higher values for both of these metrics (i.e. 28 for query similarity compared to 0.6 or lower for the others).
Table 2: Metrics measuring the density of the qrel matrix. We see that LIMIT is significantly higher than other datasets, but that the closest are instruction-following datasets such as Core17 from FollowIR. Our empirical ablations suggest (although cannot definitively prove) that datasets with higher values here will be harder for retrieval models to represent.
Dataset Name Graph Density Average Query Strength
NQ 0 0
HotPotQA 0.000037 0.1104
SciFact 0.001449 0.4222
FollowIR Core17 0.025641 0.5912
LIMIT 0.085481 28.4653
Appendix G Table Forms of Figures
In this section we show the table form of various figures. For Figure 3 it is Table 5, Figure 4 in Table 4, Figure 2 in Table 6, and Figure 5 in Table 3.
Split Dim Recall@2 Recall@10 Recall@100
Test 32 85.5 98.4 100.0
Test 64 90.4 98.7 100.0
Test 128 93.1 99.5 99.9
Test 256 94.2 99.7 100.0
Test 384 95.6 99.6 100.0
Test 512 94.0 99.5 99.9
Test 768 96.1 99.8 100.0
Test 1024 96.5 99.8 100.0
Train 32 0.0 0.0 0.0
Train 64 0.1 0.3 2.2
Train 128 0.2 0.7 3.1
Train 256 0.0 0.0 0.4
Train 384 1.1 2.7 8.3
Train 512 0.7 2.3 9.8
Train 768 0.7 2.4 9.9
Train 1024 1.0 2.8 11.2
Table 3: Fine-tuning results in table form. See Figure 5 for the comparable plot.
Model Dim Recall@2 Recall@10 Recall@20
BM25 default 97.8 100.0 100.0
E5-Mistral 7B 32 7.9 32.6 56.2
E5-Mistral 7B 64 10.2 37.0 60.3
E5-Mistral 7B 128 14.5 41.9 65.9
E5-Mistral 7B 256 15.3 45.9 69.7
E5-Mistral 7B 512 22.2 54.7 74.8
E5-Mistral 7B 768 21.6 57.5 79.2
E5-Mistral 7B 1024 24.5 60.5 80.0
E5-Mistral 7B 2048 28.9 66.3 83.2
E5-Mistral 7B 3072 29.9 67.8 85.3
E5-Mistral 7B 4096 29.5 68.1 85.2
GTE-ModernColBERT default 83.5 97.6 99.1
GritLM 7B 32 7.8 33.5 56.3
GritLM 7B 64 9.4 35.9 59.6
GritLM 7B 128 14.2 42.7 64.9
GritLM 7B 256 17.3 46.2 68.3
GritLM 7B 512 21.8 55.6 76.7
GritLM 7B 768 23.8 58.1 80.1
GritLM 7B 1024 26.2 61.4 80.1
GritLM 7B 2048 33.0 69.1 86.2
GritLM 7B 3072 36.3 72.9 89.9
GritLM 7B 4096 38.4 75.4 90.5
Promptriever Llama3 8B 32 6.1 31.4 56.0
Promptriever Llama3 8B 64 8.9 35.8 62.3
Promptriever Llama3 8B 128 13.7 44.5 67.6
Promptriever Llama3 8B 256 18.5 52.1 74.1
Promptriever Llama3 8B 512 27.0 61.8 81.7
Promptriever Llama3 8B 768 35.5 69.0 84.7
Promptriever Llama3 8B 1024 38.0 73.5 89.1
Promptriever Llama3 8B 2048 46.2 83.6 94.2
Promptriever Llama3 8B 3072 49.2 87.3 96.6
Promptriever Llama3 8B 4096 54.3 90.0 97.7
Qwen3 Embed 32 8.3 30.6 53.9
Qwen3 Embed 64 9.4 35.5 57.6
Qwen3 Embed 128 11.6 38.3 60.8
Qwen3 Embed 256 14.3 41.6 63.8
Qwen3 Embed 512 16.1 43.7 66.0
Qwen3 Embed 768 17.2 45.3 69.3
Qwen3 Embed 1024 17.8 48.7 70.3
Qwen3 Embed 2048 19.5 51.5 72.4
Qwen3 Embed 3072 19.3 52.8 73.3
Qwen3 Embed 4096 19.0 52.3 73.8
Gemini Embed 2 4.2 23.0 45.5
Gemini Embed 4 4.2 21.9 46.0
Gemini Embed 8 4.9 23.2 47.0
Gemini Embed 16 5.2 24.7 47.5
Gemini Embed 32 6.3 25.2 50.6
Gemini Embed 64 6.9 30.6 55.0
Gemini Embed 128 7.7 37.0 62.9
Gemini Embed 256 14.6 46.9 69.7
Gemini Embed 512 23.3 58.4 77.9
Gemini Embed 768 28.8 67.5 84.5
Gemini Embed 1024 31.8 69.9 86.1
Gemini Embed 2048 31.9 70.3 87.1
Gemini Embed 3072 33.7 72.4 87.9
Snowflake Arctic L 32 8.3 30.3 53.8
Snowflake Arctic L 64 9.0 35.4 58.5
Snowflake Arctic L 128 12.7 41.3 65.1
Snowflake Arctic L 256 16.0 48.2 72.6
Snowflake Arctic L 512 16.7 51.3 74.1
Snowflake Arctic L 768 17.9 53.5 74.6
Snowflake Arctic L 1024 19.4 54.9 76.0
Snowflake Arctic L 2048 19.4 54.9 76.0
Snowflake Arctic L 3072 19.4 54.9 76.0
Snowflake Arctic L 4096 19.4 54.9 76.0
Table 4: Results for the LIMIT small version. See comparable Figure 4.
Model Dim Recall@2 Recall@10 Recall@100
E5-Mistral 7B 32 0.0 0.0 0.5
E5-Mistral 7B 64 0.0 0.1 0.4
E5-Mistral 7B 128 0.1 0.3 1.0
E5-Mistral 7B 256 0.4 0.9 1.9
E5-Mistral 7B 512 0.7 1.3 3.8
E5-Mistral 7B 768 0.9 1.7 4.3
E5-Mistral 7B 1024 0.9 1.8 5.9
E5-Mistral 7B 2048 1.0 1.9 6.8
E5-Mistral 7B 3072 1.3 2.0 7.7
E5-Mistral 7B 4096 1.3 2.2 8.3
Snowflake Arctic L 32 0.0 0.1 0.6
Snowflake Arctic L 64 0.2 0.4 1.7
Snowflake Arctic L 128 0.1 0.3 1.8
Snowflake Arctic L 256 0.2 0.8 2.5
Snowflake Arctic L 512 0.3 1.0 2.5
Snowflake Arctic L 768 0.4 1.1 3.1
Snowflake Arctic L 1024 0.4 0.8 3.3
Snowflake Arctic L 2048 0.4 0.8 3.3
Snowflake Arctic L 3072 0.4 0.8 3.3
Snowflake Arctic L 4096 0.4 0.8 3.3
GritLM 7B 32 0.0 0.0 0.8
GritLM 7B 64 0.0 0.1 0.3
GritLM 7B 128 0.1 0.3 1.3
GritLM 7B 256 0.1 0.4 2.8
GritLM 7B 512 0.6 1.8 6.5
GritLM 7B 768 1.5 3.1 8.7
GritLM 7B 1024 1.8 3.5 10.6
GritLM 7B 2048 2.3 4.3 11.8
GritLM 7B 3072 2.0 4.3 12.9
GritLM 7B 4096 2.4 4.1 12.9
Promptriever Llama3 8B 32 0.0 0.0 0.1
Promptriever Llama3 8B 64 0.0 0.0 0.3
Promptriever Llama3 8B 128 0.0 0.1 0.6
Promptriever Llama3 8B 256 0.2 0.4 1.8
Promptriever Llama3 8B 512 0.6 1.4 5.4
Promptriever Llama3 8B 768 1.3 3.1 8.7
Promptriever Llama3 8B 1024 2.1 4.4 12.8
Promptriever Llama3 8B 2048 3.2 6.5 18.1
Promptriever Llama3 8B 3072 2.9 6.3 17.8
Promptriever Llama3 8B 4096 3.0 6.8 18.9
Qwen3 Embed 32 0.0 0.1 1.1
Qwen3 Embed 64 0.0 0.2 1.0
Qwen3 Embed 128 0.3 0.4 1.8
Qwen3 Embed 256 0.4 0.8 3.2
Qwen3 Embed 512 0.6 1.3 3.3
Qwen3 Embed 768 0.7 1.5 3.8
Qwen3 Embed 1024 0.7 1.6 4.6
Qwen3 Embed 2048 0.9 1.7 4.7
Qwen3 Embed 3072 0.8 1.6 4.8
Qwen3 Embed 4096 0.8 1.8 4.8
Gemini Embed 2 0.0 0.0 0.1
Gemini Embed 4 0.0 0.0 0.0
Gemini Embed 8 0.0 0.0 0.0
Gemini Embed 16 0.0 0.0 0.0
Gemini Embed 32 0.0 0.0 0.0
Gemini Embed 64 0.0 0.0 0.3
Gemini Embed 128 0.0 0.1 0.3
Gemini Embed 256 0.0 0.1 1.2
Gemini Embed 512 0.2 1.1 3.6
Gemini Embed 768 0.9 2.5 7.6
Gemini Embed 1024 1.3 2.7 8.1
Gemini Embed 2048 1.5 3.1 8.5
Gemini Embed 3072 1.6 3.5 10.0
GTE-ModernColBERT default 23.1 34.6 54.8
BM25 default 85.7 90.4 93.6
Table 5: Results on LIMIT. See comparable Figure 3.
$d$ Critical-$n$
4 10
5 14
6 19
7 24
8 28
9 32
10 36
11 42
12 47
13 54
14 62
15 70
16 79
17 89
18 99
19 109
20 120
21 132
22 144
23 157
24 170
25 184
26 198
27 213
28 229
29 245
30 261
31 278
32 296
33 314
34 333
35 352
36 372
37 392
38 413
39 434
40 460
41 484
42 505
43 545
44 605
45 626
Table 6: Critical Values of n for different d values in the Free Embedding optimization experiments. See Figure 2 for the corresponding figure.
Model BEIR LIMIT R@100
Snowflake Arctic 55.22 3.3
Promptriever 56.40 18.9
E5-Mistral 57.07 8.3
GritLM 57.40 12.9
Gemini Embed 62.65 10.0
Qwen3 Embed 62.76 4.8
Table 7: BEIR vs LIMIT results. See Figure 7 for the comparable plot.